Recent Log Entries

Follow-up Re: 2L22 Cooldown complete

Log Number: 3991807Logbook(s): Entry Makers: weaksmccryoBody: 

JT Max setpoint reduced to 60% due to overpressure causing relief valve to vent.
Lowing the JT Max allowed the relief to reseat, and allowed guard vacuum pressure to recover.

LL Setpoint reduced to 20% to allow for L-Tube - U-Tube swap to occur.

SRF Cryo procedures will be altered to include these notes, and prevent future occurrences.

COMMENT ON 1L23 3-stub tuning

Log Number: 3991806Logbook(s): Body: 

The Director's Council has decided that 1100 kW, ~500 uA max in the linac, will be the maximum beam current for the foreseeable future. From Tom's calculations an input bandwidth of 95-100 Hz is more appropriate for the lower current. No test at higher beam power is authorized. FSAD and ASE will show a paper limit of ~1200 kW so BELS can be set to provide a little headroom for small random current excursions and resulting power excursions. FSAD and ASE changes have been promised for 30 September. Maximum linac current request June-August 400 uA aka 80 uA to hall C, fifth pass.

vacuum Friday afternoon

Log Number: 3991803Logbook(s): Entry Makers: marcyBody: 

This afternoon, Leo was able to energize VIP0I04 (chopping chamber) on the gamma SPC supply in the upstairs rack, which is part of the interlock chain. Around the same time, I worked with crew chief on call to reboot iocin1 to try to restore communication to the UHV supply readings - this didn't help the UHV supplies as far as I can tell.

Removing VIP0I04 from local control and re-energizing upstairs with other supply (still recorded as VINJDIG04) caused some vacuum activity on the chopper pumps (VINJDIG07) that is recovering quickly.

IOC: in1 Crash Recovery with settings from hour normal.2021-12-21.08.49.27

Log Number: 3991802Logbook(s): Tags: Body: 

keyword=IOC: in1 Crash Recovery with settings from hour normal.2021-12-21.08.49.27
Any error messages from restore are below:

-> Preparing to restore interlock alarm parameters -- press return to continue \n
\n

-> Preparing to restore chopper parameters -- press return to continue \n

-> Preparing to restore gun parameters -- press return to continue \n

-> Preparing to restore bunch parameters -- press return to continue \n

-> Preparing to restore camacVer parameters -- press return to continue \n

-> Preparing to restore vac parameters - press return\n

-> Restoring IO

vacuum status update

Log Number: 3991801Logbook(s): Tags: Entry Makers: gramesBody: 

Vacuum progress today...

a) thanks Mark! - capture turned on this morning
b) thanks Leo! - chopping chamber returned to Dig #2 ~15:30

Fig. 1 shows vacuum levels last ~24 hours now w/ both chopper and capture skids on.

Let's see if pumps <50 uA by Monday...

iocsl24rfint restored

Log Number: 3991792Logbook(s): Tags: Body: restorelib 2.4 Target restore date was 2022-02-16 14:40:56 The output from the myRestore execution was as follows: iocsl24rfint Rf12Intlk_Analog@2022-02-16 14:40 55(55) restored Rf12Intlk_Slope@2022-02-16 14:40 121(121) restored Pause: 5 second(s) Execute: caput R2OXIARCinit 1 (success) Old : R2OXIARCinit 0 New : R2OXIARCinit 1 system@2022-02-16 14:40 2(2) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocsl24rfint (success) Procedure return status: Success

RF Module Downloads

Log Number: 3991788Logbook(s): Body: COMMENTS: Ceramic window settings downloaded.


The following module files were successfully downloaded
nl034e

RF Module Downloads

Log Number: 3991786Logbook(s): Body: COMMENTS: Ceramic window settings downloaded.


The following module files were successfully downloaded
nl038e

RF Module Downloads

Log Number: 3991785Logbook(s): Body: COMMENTS: Ceramic window settings downloaded.


The following module files were successfully downloaded
nl037e

RF Module Downloads

Log Number: 3991784Logbook(s): Body: COMMENTS: Ceramic window settings downloaded.


The following module files were successfully downloaded
nl036e

Pages